ThatCreepyBoy:

Solve 3^(2x) = 7^(x−1). −7.74293 7.74293 −1 1

ThatCreepyBoy:

(First one)

Vocaloid:

take the natural log of each side ln[3^(2x)] = ln[7^(x−1)] bring the exponents down 2x*ln(3) = (x-1)*ln(7) distribute 2x*ln(3) = x*ln(7) - ln(7) move the variables to one side 2x*ln(3) - x*ln(7) = - ln(7) factor out x x(2*ln(3) - ln(7)) = - ln(7) then isolate x and plug in the logs into your calculator to solve

ThatCreepyBoy:

Okie, one second

ThatCreepyBoy:

I got

ThatCreepyBoy:

−7.74293

Vocaloid:

good do you understand all the steps?

ThatCreepyBoy:

Yes maams :D, thank you!